i
Solve the system providing only real solutions. Use matrix methods only. X = (1 -) X LIT
Answer #1

х- Ах, then As 1. е оны ( - ) characteristic ean of A in ) = 0 2) а 8 A+ (1=0 г) а = R+2 і 4 ti 6-а - 5 2 га 2Then A KL) dik O Lee K = ( 1 ) be am eigen vector corresponding to di=(4+i). giver (6-4-1) K, - 5 K₂=0 K. +(2-4-1)ų = equival0 = e eht(cost tifint) ( 24 ) =e e46 = eut (rest i (6314 (2 cost-finf) Ii (cost t2 fint) cost t Sint 48/rcost-sint & cost t2